# To See or Not to See - Fingerprinting Devices in Adversarial Environments Amid Advanced Machine Learning

### v1

@@ Abstract
Index Terms:
Embedded Devices, Secure Communications, and Mobile and Ubiquitous Systems.
@@ I Introduction
The Internet of Things (IoT) is becoming increasingly commonplace in our lives, bringing with it numerous benefits and opportunities such as smart homes, smart manufacturing, and wearable and mobile computing. However, the integration of IoT devices in these environments and applications raises the stakes of potential security attacks. These attacks could have devastating and far-reaching consequences for affected systems, not just for the individual user but for an entire network, leading to catastrophic outcomes [1, 2, 3, 4, 5]. Therefore, it is essential to take the necessary steps to ensure that IoT systems are secure and protected from malicious actors.
Practical solutions for protecting IoT devices and networks against attacks rely on authentication. The authentication process often involves leveraging cryptographic primitives such as message authentication codes (MACs). These cryptographic-based solutions provide provable security guarantees, however, they impose large overheads on the IoT system. This is an important concern since IoT devices are typically resource-constrained, and in many cases, solutions with large power and/or storage overhead are not practical and/or feasible. Moreover, cryptographic-based methods are subject to various side-channel and/or physical attacks [6, 7, 8, 9, 10].
@@
With the recent strides in machine learning, particularly in generative models like GANs, LLMs, and diffusion models [11, 12, 13, 14, 15, 16, 17], it is essential to extensively examine the strength of current solutions against these types of attacks. This paper aims to expand on prior fingerprinting research surveys [18, 19, 20, 21], delving into a systematic exploration of attack and defense trends for IoT device fingerprinting and eavesdropping. Furthermore, it delves into future attack trends in light of the advancements in machine learning.
@@ III Device Authentication
RF WiFi Packet Averaging, Filtering, Slicing, Partial Equalization CFO, I/Q Offset/Imbalance, Amplitudes Stats., F.B., D.L. [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]
BLE Packet Estimation, GATT CFO and IQ Offset, GATT Stats. [2, 33]
ZigBee Packet Synchronization, Offset Calculation, DRA DCTF, CFO, ROI F.B., D.L. [34, 35, 36, 37, 38]
LoRa Packet Decoding, Synchronization, Normalization Tx Imperfections, IQ Features F.B., D.L. [39, 40, 41, 42]
RFID RFID Signal Resolution Alignment, Sequencing, EPC Identification Persistence Time, Normalization, Tag Imperfection, Peak Frequency Stats., F.B., D.L. [43, 44, 1]
QPSK Symbols DFT, Averaging Power Amplifier Nonlinearity D.L. [45]
WiMAX Near-Transient Response Filtering Standard Deviation, Variance, Skewness, Kurtosis F.B. [46]
Drones I/Q Samples Noise Reduction, Trace Segmentation Parasitic Response, D.L. [47, 48]
Network Packet Flows Traffic Parsing and Collection of Data Timing, Clock Offsets, RSS, Flow Volume Stats., F.B. [49, 50, 51, 52, 53, 18]
EM I/Q Samples Averaging, Folding, FFT, STFT, Normalization Clock, IC signature, Magnetic Induction, Ambient EM Stats., F.B., D.L. [54, 55, 56, 57, 58, 59, 60, 61, 62]
Range - $<$ 10 m,   10-50 m, $>$ 50 m
@@ III-A RF Fingerprinting
Input. For RF Fingerprinting systems, data can be received as an RF signal [44, 1, 43], I/Q samples [39, 41, 24, 48, 25, 35, 27], and sequences of packets [22, 23]. In many cases, a preprocessing step is required to process and prepare the data such that it is in a proper format for feature extraction and classification.
Preprocessing and Features. When an RF signal is transmitted, the signal travels through a communication channel, which may have a negative effect (such as distortion) on the signal due to differences in the observed channel impulse response and additive noise. Therefore, various signal processing methods such as filtering can be applied to eliminate these effects  [39, 32, 23, 26, 24] as well as reduce the effects of noise  [47, 54, 26] or device movement [47].
The signal can be processed in different domains such as in the time domain, frequency domain, or both by using a Short-Time Fourier Transform (STFT)  [40]. In short, different signal processing methods can be employed to transform a received signal to the desired format.
@@
Common features used in prior work are shown in Table I. As depicted in the table, there are many types of features such as peak frequency, frequency offset, signal-to-noise ratio (SNR), and statistical features [30, 2]. Hardware imperfections also impart unique features onto the sent or measured signal which allows us to extract unique features for fingerprinting  [32, 25, 27, 42, 45, 47].
Classifier. Many approaches are used in the RF space for classification. Deep learning is a prevalent approach for many modalities such as WiFi [23, 26, 27, 24, 25], ZigBee [36] and LoRa [42]. Feature-based approaches such as support vector machine and kNN models are also used [29, 46, 1]. Statistical approaches are also commonly used such as distance-based metrics  [2], profile analysis  [33], and statistical t-tests  [44].
Performance. The detection range of RF fingerprinting approaches often depends on the modality, but was generally short. For instance, LoRa typically has a range in the thousands of meters but was tested in a 2240 square feet room in DeepLoRa [40]. RFID, a short-range modality, understandably has a short range. Interestingly, in a work that fingerprinted drones, the range was short (around 8 meters), due to the distance limitations of mm-Wave signals  [47].
The scalability of the systems varies as well. WiFi, BLE, and RFID are the most scalable, in that there are works that test thousands of devices [26, 33, 44]. However, the number of devices is typically tens to hundreds. One LoRa work tested 60 LoRa devices [39], and many WiFi works tested in the hundreds of devices [24, 28, 29]. Spoofing is considered in a few works, but not in many. Rogue devices are considered in a few LoRa and ZigBee works [38, 39].
@@ III-B Network Fingerprinting
Preprocessing and Features. In network approaches, typical preprocessing strategies include traffic parsing [49] and the collection of data [51, 53]. Once data is preprocessed, we then isolate specific features, such as the timing of packets [49], clock offsets [50], RSS [18], and flow volume [53].
Classifier. In network approaches, we typically see statistically-based or feature-based classifiers. Similarity metrics are often used as classifier models, such as in IoTFinder and many others  [49, 52, 51]. We see the usage of feature-based classifiers such as in one work that utilizes Naive Bayes and a random forest classifier [53].
@@
Network fingerprinting approaches generally consider devices in the region of tens of devices. Two approaches consider device counts in the twenties [52, 53]. Two other approaches considered 53 devices [49] and 66 idle devices [51].
Most works do not consider the impacts of spoofing and adversarial attacks. However, a select few consider adversaries and rogue APs  [52, 18].
@@ III-C Electromagnetic Fingerprinting
Electromagnetic fingerprinting relies on measuring electromagnetic signals (EM) emitted from processors, memory, or hardware units. Oftentimes these signals are produced unintentionally via programmatic activity and have typically been used as a side-channel [10, 63, 64]. For fingerprinting, properties of these EM signals caused by uniqueness in the hardware itself can be used for authentication.
Input. In EM fingerprinting, inputs are usually EM signals that are emitted from different electronic devices, CPU, and memory activities. On the receiver end, the information is typically received as I/Q samples [60, 62].
Preprocessing and Features. Like other fingerprinting works, preprocessing is critical towards extracting features. Methods such as averaging  [54] and filtering  [60, 62] are typically used to reduce noise and prepare the signal for analysis [54].
The next step after performing preprocessing is understanding the characteristics of the devices we are authenticating. Different pieces of hardware such as the CPU  [55, 62], DRAM  [60], and integrated circuits  [59, 58] emanate unique electromagnetic emanations per device. Generally speaking, electromagnetic fingerprinting approaches utilize the unique frequency response of circuits, processors, and memory components to identify between devices.
Classifier. In EM Fingerprinting, a wide range of classifiers are used. Statistical approaches like cosine similarity are used in EM-ID and other approaches [57, 58, 59]. Additionally, like RF fingerprinting, feature-based methods are a common strategy  [60, 55, 62].
Performance. Range typically falls short in electromagnetic approaches. Many methods utilize an electromagnetic probe to measure the signal, resulting in near-field ranges [57, 55, 58, 59]. Digitus  [62] and MemScope  [60] are exceptions with a range of around 10 meters and 30 meters, respectively, via leveraging CPU and memory clock signals.
Scalability is typically in the tens of devices. On the short side, one approach only tests 5 same type devices [56]. EM-ID tests 40 devices [57], MemScope tests 65 devices [60], and DeMiCPU tests 90 devices (70 laptops and 20 phones) [55].
Unfortunately, outside of Digitus and MemScope, none of the papers reviewed discussed spoofing or adversarial attacks in detail. Digitus considers spoofing and hidden devices  [62], whereas MemScope discusses the classification of alien devices as well as record-and-replay attacks [60].
{mdframed}
[backgroundcolor=Apricot]
@@ IV Eavesdropping Detection
Lumos [65] Identify hidden WiFi-connected devices in a room via WiFi packets. RF Packet Packet Features F.B. Cameras, mics, doorbells, etc. Macbook or RPI
Ghostbuster [4] Detect wireless eavesdroppers (passive receivers) via unintended RF leakage. RF Packet Passive RF Leakage Signal Processing USRP USRP
E-Eye [66] Find hidden e-devices via mm-Wave sensing. RF mm-Wave Signal Non Linear Effects F.B. Laptop, tablet, wearable, etc. mm-Wave Probe
Blink + Flicker [67] Detect spy cameras via changing ambient light and analyzing packet flow changes. Network Packet Flows Packet Flow Changes Stats. Wireless IP Cameras Linux Tablet / RPI
SnoopDog [3] Discovering and localizing wireless sensors via user action and wireless traffic. Network Packet Flows Traffic Patterns Database Wireless Sensors Mobile Phone
EarFisher [68] Find eavesdroppers by analyzing their inherent memory EMR response. EM I/Q Samples Memory EMR Stats. N/A Laptop BladeRF
HeatDeCam [69] Detect hidden spy cameras without wireless connectivity using thermal imagery. Thermal Thermal Input, Image Heat Dissipation D.L. Spy Cameras Mobile Phone
Range - $<$ 1 m, $<$ 5 m, $>$ 5 m
Time - $<$ 1 min, $<$ 5 min, $>$ 5 min
@@ IV-A RF Approaches
RF hidden device detection attempts to discover a hidden device via either their packet features [65] or a device’s unique response [4, 66]. In the first case, we need to listen to the normal activity of the device. In the second case, by sending a specific input, we can elicit a certain detectable response from the device in order to detect it. In the following, more details about the various steps in this process are provided.
Input. For RF devices, hidden device detection depends on the transmission of packets via transmitter modules and the resulting packets or I/Q samples that the detector receives. For some approaches, the packets sent from the hidden device are used to detect the device itself. Many works utilize WiFi packets in their detection system [65, 4, 70, 71]. Other works depend on a stimulus or an input that invokes a response from the device. These approaches utilize the raw I/Q samples from the device [66, 72, 73, 74, 75]. In either case, the detectors receive the data as input, conduct any preprocessing (averaging, smoothing, etc.) needed to prepare the signal for classification, then send the preprocessed data to the detector.
Features. The features used to detect a device depend on the received input. In many cases, receiver leakage or characteristics of unintended emissions can be used as features to detect a device [66, 4, 72, 73, 74, 75, 71]. In other cases, features from the packet itself, such as RSS information, can be used [65, 70].
Classifiers. How the above features are used depends on the classifier involved in the detection. Various kinds of classifiers are used to detect the presence of hidden devices, ranging from lightweight statistical approaches to complex machine learning approaches. Many approaches utilize signal processing and statistics to identify hidden devices  [71, 72]. Some common techniques include filtering, sampling, applying transformations, and subtracting out unwanted signals  [4, 73, 74].
Some approaches utilize a feature-based approach. A feature-based approach depends on identifying a unique response from a device. This unique response could be features analyzed from a packet  [65], or from features derived from device measurements  [66]. These features can be then used as inputs into lightweight models such as XGBoost and support vector machine  [65, 66]. Beyond these approaches, one can utilize RSS as input for localization algorithms [70]. RSS information can be received within the packet and be used as input to a log-distance path loss model for instance.
@@
For the receiving device, a common device used is a software-defined radio [4, 73, 74, 71]. Laptops or smartphones are often used as well [65, 70]. These devices are often used for wireless reception.
A variety of hidden devices are tested including other laptops [66], cameras [65, 70], microphones [65], and GMRS radios [73, 74].
The range of detection varies depending on the technique and technology. Detection can be near-field in mm-Wave applications  [66] or within the space of a room such as in Lumos [65] and Ghostbuster [4]. Finally, approaches such as Matthan can achieve tens to hundreds of meters [74, 71].
The time of detection varies as well. Some detect under a minute [4, 75] while others take many minutes [65, 70]. A shorter detection time is better because it reduces the time the user could potentially be exposed to an adversary.
Finally, only a few approaches localize the device. In Lumos [65], the device is localized within 1.5 meters. Two other approaches localize the device under 5 meters [74, 70].

### v2

@@ I Introduction
The Internet of Things (IoT) is becoming increasingly commonplace in our lives, bringing with it numerous benefits and opportunities such as smart homes, smart manufacturing, and wearable and mobile computing. However, the integration of IoT devices in these environments and applications raises the stakes of potential security attacks. These attacks could have devastating and far-reaching consequences for affected systems, not just for the individual user but for an entire network, leading to catastrophic outcomes [joo2020hold, givehchian2022evaluating, singh2021always, chaman2018ghostbuster, bao2021threat]. Therefore, it is essential to take the necessary steps to ensure that IoT systems are secure and protected from malicious actors.
Practical solutions for protecting IoT devices and networks against attacks rely on authentication. The authentication process often involves leveraging cryptographic primitives such as message authentication codes (MACs). These cryptographic-based solutions provide provable security guarantees, however, they impose large overheads on the IoT system. This is an important concern since IoT devices are typically resource-constrained, and in many cases, solutions with large power and/or storage overhead are not practical and/or feasible. Moreover, cryptographic-based methods are subject to various side-channel and/or physical attacks [zhao2018fpga, genkin2015get, lipp2021platypus, guri2015gsmem, camurati2020understanding].
@@
With the recent strides in machine learning, particularly in generative models like GANs, LLMs, and diffusion models [li2021gnss, ayanoglu2022machine, creswell2018generative, chen2021evaluating, wei2022emergent, croitoru2023diffusion, yang2022diffusion], it is essential to extensively examine the strength of current solutions against these types of attacks. This paper aims to expand on prior fingerprinting research surveys [xu2015device, soltanieh2020review, liu2021machine, xie2020survey], delving into a systematic exploration of attack and defense trends for IoT device fingerprinting and eavesdropping. Furthermore, it delves into future attack trends in light of the advancements in machine learning.
@@ III Device Authentication
RF WiFi Packet Averaging, Filtering, Slicing, Partial Equalization CFO, I/Q Offset/Imbalance, Amplitudes Stats., F.B., D.L. [yu2020you, hanna2020open, jian2021radio, al2020exposing, jian2020deep, sankhe2019no, restuccia2019deepradioid, brik2008wireless, hua2018accurate, azarmehr2017wireless, agadakos2020chameleons]
BLE Packet Estimation, GATT CFO and IQ Offset, GATT Stats. [givehchian2022evaluating, celosia2019fingerprinting]
ZigBee Packet Synchronization, Offset Calculation, DRA DCTF, CFO, ROI F.B., D.L. [peng2018design, merchant2018deep, yu2019robust, merchant2019enhanced, bihl2016feature]
LoRa Packet Decoding, Synchronization, Normalization Tx Imperfections, IQ Features F.B., D.L. [shen2022towards, al2021deeplora, zhang2021radio, das2018deep]
RFID RFID Signal Resolution Alignment, Sequencing, EPC Identification Persistence Time, Normalization, Tag Imperfection, Peak Frequency Stats., F.B., D.L. [pan2022rf, chen2020eingerprint, joo2020hold]
QPSK Symbols DFT, Averaging Power Amplifier Nonlinearity D.L. [hanna2019deep]
WiMAX Near-Transient Response Filtering Standard Deviation, Variance, Skewness, Kurtosis F.B. [reising2020radio]
Drones I/Q Samples Noise Reduction, Trace Segmentation Parasitic Response, D.L. [li2022reliable, soltani2020rf]
Network Packet Flows Traffic Parsing and Collection of Data Timing, Clock Offsets, RSS, Flow Volume Stats., F.B. [perdisci2020iotfinder, kohno2005remote, charyyev2020locality, miettinen2017iot, sivanathan2018classifying, xu2015device]
EM I/Q Samples Averaging, Folding, FFT, STFT, Normalization Clock, IC signature, Magnetic Induction, Ambient EM Stats., F.B., D.L. [lee2022aerokey, cheng2019demicpu, mariano2019classification, yang2016id, ahmed2018authentication, ahmed2017radiated, shen2022electromagnetic, ibrahim2020magneto, feng2023fingerprinting]
Range - << 10 m,   10-50 m, >> 50 m
@@ III-A RF Fingerprinting
Input. For RF Fingerprinting systems, data can be received as an RF signal [chen2020eingerprint, joo2020hold, pan2022rf], I/Q samples [shen2022towards, zhang2021radio, jian2021radio, soltani2020rf, al2020exposing, merchant2018deep, sankhe2019no], and sequences of packets [yu2020you, hanna2020open]. In many cases, a preprocessing step is required to process and prepare the data such that it is in a proper format for feature extraction and classification.
Preprocessing and Features. When an RF signal is transmitted, the signal travels through a communication channel, which may have a negative effect (such as distortion) on the signal due to differences in the observed channel impulse response and additive noise. Therefore, various signal processing methods such as filtering can be applied to eliminate these effects  [shen2022towards, agadakos2020chameleons, hanna2020open, jian2020deep, jian2021radio] as well as reduce the effects of noise  [li2022reliable, lee2022aerokey, jian2020deep] or device movement [li2022reliable].
The signal can be processed in different domains such as in the time domain, frequency domain, or both by using a Short-Time Fourier Transform (STFT)  [al2021deeplora]. In short, different signal processing methods can be employed to transform a received signal to the desired format.
@@
Common features used in prior work are shown in Table I. As depicted in the table, there are many types of features such as peak frequency, frequency offset, signal-to-noise ratio (SNR), and statistical features [hua2018accurate, givehchian2022evaluating]. Hardware imperfections also impart unique features onto the sent or measured signal which allows us to extract unique features for fingerprinting  [agadakos2020chameleons, al2020exposing, sankhe2019no, das2018deep, hanna2019deep, li2022reliable].
Classifier. Many approaches are used in the RF space for classification. Deep learning is a prevalent approach for many modalities such as WiFi [hanna2020open, jian2020deep, sankhe2019no, jian2021radio, al2020exposing], ZigBee [yu2019robust] and LoRa [das2018deep]. Feature-based approaches such as support vector machine and kNN models are also used [brik2008wireless, reising2020radio, joo2020hold]. Statistical approaches are also commonly used such as distance-based metrics  [givehchian2022evaluating], profile analysis  [celosia2019fingerprinting], and statistical t-tests  [chen2020eingerprint].
Performance. The detection range of RF fingerprinting approaches often depends on the modality, but was generally short. For instance, LoRa typically has a range in the thousands of meters but was tested in a 2240 square feet room in DeepLoRa [al2021deeplora]. RFID, a short-range modality, understandably has a short range. Interestingly, in a work that fingerprinted drones, the range was short (around 8 meters), due to the distance limitations of mm-Wave signals  [li2022reliable].
The scalability of the systems varies as well. WiFi, BLE, and RFID are the most scalable, in that there are works that test thousands of devices [jian2020deep, celosia2019fingerprinting, chen2020eingerprint]. However, the number of devices is typically tens to hundreds. One LoRa work tested 60 LoRa devices [shen2022towards], and many WiFi works tested in the hundreds of devices [jian2021radio, restuccia2019deepradioid, brik2008wireless]. Spoofing is considered in a few works, but not in many. Rogue devices are considered in a few LoRa and ZigBee works [bihl2016feature, shen2022towards].
@@ III-B Network Fingerprinting
Preprocessing and Features. In network approaches, typical preprocessing strategies include traffic parsing [perdisci2020iotfinder] and the collection of data [charyyev2020locality, sivanathan2018classifying]. Once data is preprocessed, we then isolate specific features, such as the timing of packets [perdisci2020iotfinder], clock offsets [kohno2005remote], RSS [xu2015device], and flow volume [sivanathan2018classifying].
Classifier. In network approaches, we typically see statistically-based or feature-based classifiers. Similarity metrics are often used as classifier models, such as in IoTFinder and many others  [perdisci2020iotfinder, miettinen2017iot, charyyev2020locality]. We see the usage of feature-based classifiers such as in one work that utilizes Naive Bayes and a random forest classifier [sivanathan2018classifying].
@@
Network fingerprinting approaches generally consider devices in the region of tens of devices. Two approaches consider device counts in the twenties [miettinen2017iot, sivanathan2018classifying]. Two other approaches considered 53 devices [perdisci2020iotfinder] and 66 idle devices [charyyev2020locality].
Most works do not consider the impacts of spoofing and adversarial attacks. However, a select few consider adversaries and rogue APs  [miettinen2017iot, xu2015device].
@@ III-C Electromagnetic Fingerprinting
Electromagnetic fingerprinting relies on measuring electromagnetic signals (EM) emitted from processors, memory, or hardware units. Oftentimes these signals are produced unintentionally via programmatic activity and have typically been used as a side-channel [camurati2020understanding, nazari2017eddie, he2017hardware]. For fingerprinting, properties of these EM signals caused by uniqueness in the hardware itself can be used for authentication.
Input. In EM fingerprinting, inputs are usually EM signals that are emitted from different electronic devices, CPU, and memory activities. On the receiver end, the information is typically received as I/Q samples [shen2022electromagnetic, feng2023fingerprinting].
Preprocessing and Features. Like other fingerprinting works, preprocessing is critical towards extracting features. Methods such as averaging  [lee2022aerokey] and filtering  [shen2022electromagnetic, feng2023fingerprinting] are typically used to reduce noise and prepare the signal for analysis [lee2022aerokey].
The next step after performing preprocessing is understanding the characteristics of the devices we are authenticating. Different pieces of hardware such as the CPU  [cheng2019demicpu, feng2023fingerprinting], DRAM  [shen2022electromagnetic], and integrated circuits  [ahmed2017radiated, ahmed2018authentication] emanate unique electromagnetic emanations per device. Generally speaking, electromagnetic fingerprinting approaches utilize the unique frequency response of circuits, processors, and memory components to identify between devices.
Classifier. In EM Fingerprinting, a wide range of classifiers are used. Statistical approaches like cosine similarity are used in EM-ID and other approaches [yang2016id, ahmed2018authentication, ahmed2017radiated]. Additionally, like RF fingerprinting, feature-based methods are a common strategy  [shen2022electromagnetic, cheng2019demicpu, feng2023fingerprinting].
Performance. Range typically falls short in electromagnetic approaches. Many methods utilize an electromagnetic probe to measure the signal, resulting in near-field ranges [yang2016id, cheng2019demicpu, ahmed2018authentication, ahmed2017radiated]. Digitus  [feng2023fingerprinting] and MemScope  [shen2022electromagnetic] are exceptions with a range of around 10 meters and 30 meters, respectively, via leveraging CPU and memory clock signals.
Scalability is typically in the tens of devices. On the short side, one approach only tests 5 same type devices [mariano2019classification]. EM-ID tests 40 devices [yang2016id], MemScope tests 65 devices [shen2022electromagnetic], and DeMiCPU tests 90 devices (70 laptops and 20 phones) [cheng2019demicpu].
Unfortunately, outside of Digitus and MemScope, none of the papers reviewed discussed spoofing or adversarial attacks in detail. Digitus considers spoofing and hidden devices  [feng2023fingerprinting], whereas MemScope discusses the classification of alien devices as well as record-and-replay attacks [shen2022electromagnetic].
@@ IV Eavesdropping Detection
Lumos [sharma2022lumos] Identify hidden WiFi-connected devices in a room via WiFi packets. RF Packet Packet Features F.B. Cameras, mics, doorbells, etc. Macbook or RPI
Ghostbuster [chaman2018ghostbuster] Detect wireless eavesdroppers (passive receivers) via unintended RF leakage. RF Packet Passive RF Leakage Signal Processing USRP USRP
E-Eye [li2018eye] Find hidden e-devices via mm-Wave sensing. RF mm-Wave Signal Non Linear Effects F.B. Laptop, tablet, wearable, etc. mm-Wave Probe
Blink + Flicker [liu2018detecting] Detect spy cameras via changing ambient light and analyzing packet flow changes. Network Packet Flows Packet Flow Changes Stats. Wireless IP Cameras Linux Tablet / RPI
SnoopDog [singh2021always] Discovering and localizing wireless sensors via user action and wireless traffic. Network Packet Flows Traffic Patterns Database Wireless Sensors Mobile Phone
EarFisher [shen2021earfisher] Find eavesdroppers by analyzing their inherent memory EMR response. EM I/Q Samples Memory EMR Stats. N/A Laptop BladeRF
HeatDeCam [yu2022heatdecam] Detect hidden spy cameras without wireless connectivity using thermal imagery. Thermal Thermal Input, Image Heat Dissipation D.L. Spy Cameras Mobile Phone
Range - << 1 m, << 5 m, >> 5 m
Time - << 1 min, << 5 min, >> 5 min
@@ IV-A RF Approaches
RF hidden device detection attempts to discover a hidden device via either their packet features [sharma2022lumos] or a device’s unique response [chaman2018ghostbuster, li2018eye]. In the first case, we need to listen to the normal activity of the device. In the second case, by sending a specific input, we can elicit a certain detectable response from the device in order to detect it. In the following, more details about the various steps in this process are provided.
Input. For RF devices, hidden device detection depends on the transmission of packets via transmitter modules and the resulting packets or I/Q samples that the detector receives. For some approaches, the packets sent from the hidden device are used to detect the device itself. Many works utilize WiFi packets in their detection system [sharma2022lumos, chaman2018ghostbuster, li2018adversarial, nguyen2017matthan]. Other works depend on a stimulus or an input that invokes a response from the device. These approaches utilize the raw I/Q samples from the device [li2018eye, park2006hidden, stagner2011practical, stagner2012locating, wild2005detecting]. In either case, the detectors receive the data as input, conduct any preprocessing (averaging, smoothing, etc.) needed to prepare the signal for classification, then send the preprocessed data to the detector.
Features. The features used to detect a device depend on the received input. In many cases, receiver leakage or characteristics of unintended emissions can be used as features to detect a device [li2018eye, chaman2018ghostbuster, park2006hidden, stagner2011practical, stagner2012locating, wild2005detecting, nguyen2017matthan]. In other cases, features from the packet itself, such as RSS information, can be used [sharma2022lumos, li2018adversarial].
Classifiers. How the above features are used depends on the classifier involved in the detection. Various kinds of classifiers are used to detect the presence of hidden devices, ranging from lightweight statistical approaches to complex machine learning approaches. Many approaches utilize signal processing and statistics to identify hidden devices  [nguyen2017matthan, park2006hidden]. Some common techniques include filtering, sampling, applying transformations, and subtracting out unwanted signals  [chaman2018ghostbuster, stagner2011practical, stagner2012locating].
Some approaches utilize a feature-based approach. A feature-based approach depends on identifying a unique response from a device. This unique response could be features analyzed from a packet  [sharma2022lumos], or from features derived from device measurements  [li2018eye]. These features can be then used as inputs into lightweight models such as XGBoost and support vector machine  [sharma2022lumos, li2018eye]. Beyond these approaches, one can utilize RSS as input for localization algorithms [li2018adversarial]. RSS information can be received within the packet and be used as input to a log-distance path loss model for instance.
@@
For the receiving device, a common device used is a software-defined radio [chaman2018ghostbuster, stagner2011practical, stagner2012locating, nguyen2017matthan]. Laptops or smartphones are often used as well [sharma2022lumos, li2018adversarial]. These devices are often used for wireless reception.
A variety of hidden devices are tested including other laptops [li2018eye], cameras [sharma2022lumos, li2018adversarial], microphones [sharma2022lumos], and GMRS radios [stagner2011practical, stagner2012locating].
The range of detection varies depending on the technique and technology. Detection can be near-field in mm-Wave applications  [li2018eye] or within the space of a room such as in Lumos [sharma2022lumos] and Ghostbuster [chaman2018ghostbuster]. Finally, approaches such as Matthan can achieve tens to hundreds of meters [stagner2012locating, nguyen2017matthan].
The time of detection varies as well. Some detect under a minute [chaman2018ghostbuster, wild2005detecting] while others take many minutes [sharma2022lumos, li2018adversarial]. A shorter detection time is better because it reduces the time the user could potentially be exposed to an adversary.
Finally, only a few approaches localize the device. In Lumos [sharma2022lumos], the device is localized within 1.5 meters. Two other approaches localize the device under 5 meters [stagner2012locating, li2018adversarial].